# Wet Hair: Global Symmetries in Entanglement Islands

### v1

@@ Abstract
A central conjecture in quantum gravity is the non-existence of global symmetries. As a fully unitary theory, there is no information loss in a UV complete quantum gravity theory. We see both these concepts reflected in the AdS/CFT correspondence, which tells us that dynamical processes in AdS are fully captured by a manifestly unitary CFT with no information loss. Furthermore, global symmetries of the CFT are dual to gauge symmetries in the AdS, which implies no global symmetry in the AdS. In this work, we provide concrete evidence for the connection between the non-existence of global symmetries and the absence of information loss in quantum gravity. We study the island setups in which a gravitational AdS is coupled with a nongravitational bath on its boundary. In such theories, the information in the AdS can be lost to the bath. We provide concrete examples with global symmetries in the island setup, from both the bottom-up and the top-down perspectives. We argue that these global symmetries are consistent due to entanglement islands, in which holography is realized in a novel fashion. The global symmetries we construct are all mixed with spontaneously broken gauge symmetries. We will show that this fact has two implications: 1) The black hole hair is detectable in the bath (“wet hair”); 2) a resolution of a puzzle proposed by Harlow and Shagoulian.
@@ 1 Introduction
It has long been conjectured that global symmetries cannot exist in a UV complete quantum gravity theory Banks:1988yz ; Giddings:1988cx ; Dine:1992vx ; Kallosh:1995hi ; ArkaniHamed:2006dz ; Banks:2010zn and therefore symmetries should be either gauged or broken. This conjecture about the absence of global symmetry applies for both continuous and discrete global symmetries. In this paper, we focus on continuous global symmetries. The basic argument for the absence of this type of global symmetries makes use of the Bekenstein-Hawking entropy formula for a black hole
@@
and the no-hair theorem of black holes. The Bekenstein-Hawking formula suggests that a black hole has a finite number of microstates upper bounded by $e^{S_{BH}}$ and the no-hair theorem states that black hole solutions are independent of the global symmetry charge inside the black hole. The latter suggests that coarse-grained properties of black holes are not affected by adding any amount of global symmetry charges, which however is in tension with the former as it suggests that there would exist an infinite number of black hole microstates distinguished by their global symmetry charges. Hence, global symmetries had better not exist in quantum gravity. We note that within AdS/CFT, a putative proof of the absence of global symmetries has been put forward in Harlow:2018jwu ; Harlow:2018tng (see also Chen:2020ojn ; Hsin:2020mfa ; Bah:2022uyz for some other attempts from a rather different perspective).
@@
The finite Hilbert space the logarithm of whose dimension scales linearly with the black hole horizon area is the early motivation for the holographic principle tHooft:1993dmi ; Susskind:1994vu . This principle is explicitly realized by the AdS/CFT correspondence Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj , which states that quantum gravitational dynamics in an asymptotically anti-de Sitter (AdS) spacetime is dual to a conformal field theory (CFT) living on its asymptotic boundary. In the context of a black hole in AdS space, the quantum mechanical system that describes its physics is the dual boundary CFT, which is manifestly unitary. Hence, the Bekenstein-Hawking entropy formula suggests that the black hole dynamics, or the quantum gravitational theory that describes the black hole, is unitary. We also have seen from the previous paragraph that the Bekenstein-Hawking entropy formula suggests the absence of global symmetries in quantum gravity. Therefore, it is natural to conjecture that the absence of global symmetry is tied to the unitarity of quantum gravity.
@@
In this paper, we provide concrete evidence for the above conjecture from a novel perspective. We study a system in which a gravitational asymptotically AdS spacetime is coupled to a nongravitational bath on its asymptotic boundary, so that the gravitational theory is no longer unitary.  [^1] In principle, one can violate unitarity for a finite time, in which case our arguments still apply but the global symmetry can exist only increasingly deeper in the bulk. Such setups, with a coupling to a bath, contain entanglement islands, which have been important to recent developments in quantum gravity Almheiri:2019psf ; Penington:2019npb ; Geng:2020qvw ; Chen:2020uac ; Chen:2020hmv , so we call them island setups for convenience. Notice that our definition of island setups does not include all theories with islands, but only those in which unitarity is violated by the coupling of the gravitational system to a bath. The so-called Page curve of the black hole radiation can be computed in this setup, which suggests that the dynamics of the union of the gravitational AdS and the bath is unitary. However, in this system the dynamics in the gravitational AdS alone is not unitary due to the bath coupling. Thus, if global symmetries exist in such systems, it would demonstrate the connection between unitarity and the absence of global symmetries. More precisely, one should expect that particles charged under global symmetries can propagate into the gravitational AdS from the bath. Nevertheless, one might naively think that this expectation violates our former considerations of black holes in AdS. We will show in this paper that this is not the case and the reason is the existence of entanglement islands. We will show how entanglement islands allow for consistency of global symmetries with the above considerations of black holes.
@@
We will furthermore provide explicit examples of global symmetries in the island setup from both bottom-up holographic constructions and top-down string theory constructions. With the general argument that global symmetries can exist in the island setup and various concrete examples, we prove the early conjecture that global symmetries can exist in quantum gravitational systems with disordering Benini:2022hzx . Moreover, our work provides firm evidence for the conjecture that unitarity is essential to the absence of global symmetries in quantum gravity. To be clear we are considering only theories in which the full system preserves unitarity, but the AdS system on its own does not.
@@
We note that Ref. Harlow:2020bee proposed that unitarity and the absence of global symmetries in gravitational theories can be connected. We attack this question with different considerations and conclusions in specific examples. We should also mention the possibility that this connection might be violated in some low dimensional theories Harlow:2020bee . We do not address this question here but focus on the role of unitarity in quantum gravity theories which have propagating gravitational degrees of freedom.
@@
This paper is organized as follows. In Sec. 2 we review the physics of entanglement islands and provide a general argument of why global symmetry can exist in the island setup. In Sec. 3 we provide an explicit construction of a global symmetry in the island setup. In Sec. 4 we study the holographic dual of the island setup, i.e. the Karch-Randall braneworld Karch:2000ct ; Karch:2000gx and we construct global symmetries in the dual island setup using holography. In Sec. 5 we provide constructions of global symmetries for the dual island setup in the string theory realization of the Karch-Randall braneworld. In Sec. 6 we resolve a puzzle proposed by Harlow and Shagoulian in Harlow:2020bee where they argued that global symmetries are in tension with entanglement islands. We conclude our paper with discussions in Sec. 7.
@@ 2.1 A Lightening Review of Entanglement Islands
Entanglement islands emerged from attempts to compute the Page curve of the radiation from a black hole. This was a hard question for both technical and conceptual reasons. At the conceptual level, an evaporating black hole’s radiation and the black hole itself are living in the same spacetime, so it is unclear how to isolate the radiation from the black hole. At the technical level, it is unclear how the computation of fine-grained properties like the entanglement entropy can be done in a gravitational spacetime. Thus, a natural setup that can circumvent the above difficulties is one with a black hole spacetime coupled to a non-gravitational bath by a transparent boundary condition for energy on its boundary. Black holes in AdS have a nice asymptotic structure which allows the transparent boundary condition to be clearly modeled Porrati:2001gx ; Porrati:2002dt ; Porrati:2003sa ; Duff:2004wh ; Aharony:2006hz ; Geng:2023ynk (see Fig.1). In this setup, one can think of the non-gravitational bath as extracting the radiation from the black hole, so one can study the black hole radiation by studying this non-gravitational bath (see Fig.2).
An interesting discovery from studying the above setup is the entanglement island. It comes from the attempt to compute the entanglement entropy of the black hole radiation. The black hole radiation can be thought of as a subregion $R$ in the non-gravitational bath. The entanglement entropy of it can be computed using the quantum extremal surface formula Engelhardt:2014gca
@@
where $\mathcal{I}$ is a subregion in the gravitational AdS black hole spacetime, $S_{\text{QFT}}(R\cup\mathcal{I})$ is the entanglement entropy for the quantum fields in the subregion $R\cup\mathcal{I}$ as if it is in a fixed non-gravitational background, $A(\partial\mathcal{I})$ is the area of the boundary of the subregion $\mathcal{I}$ and $G_{N}$ is the AdS Newton’s constant. In the above formula, the subregion $\mathcal{I}$ is called the entanglement island of $R$. This formula can be derived by using holography to translate the field theory replica path integral for $S(R)$ to a gravitational path integral in the full island setup and the replica wormhole saddle in the resulting gravitational path integral proves Equ. (2) Almheiri:2019qdq ; Penington:2019kki ; Geng:2024xpj ; Geng:2025efs . The results of the formula Equ. (2) are the value of the entanglement entropy of the bath subregion $R$ and the location of its entanglement island $\mathcal{I}$ within the gravitational AdS black hole spacetime. The dominance of the entanglement island in the path integral emerges at late time (see Fig.3) and the resulting $S(R)$ obeys a unitary Page curve (see Fig.4).
Physically, one can think that the bath subregion $R$ is capturing more and more radiation from the black hole as time evolves (see Fig.2). Thus, its entanglement entropy is initially growing with time according to Hawking’s calculation Hawking:1974sw and this growth is saying that the entanglement between the black hole and the radiation it already emitted is becoming stronger. Nevertheless, a basic consistency from quantum mechanics is that the entanglement entropy between two quantum systems, i.e. the black hole and the radiation $R$, is upper bounded by the smallest Hilbert space dimension among these two quantum systems, i.e.
@@
However, a careful interpretation is required if one compares the above results with Hawking’s calculation Hawking:1974sw . Hawking’s calculation can be interpreted heuristically as virtual pairs of particles being continuously produced near the black hole horizon, due to the lack of a global time translation invariance extending into that region, and each pair consists of two particles with one outside black hole horizon and one inside (see Fig.5). The two particles in each pair are maximally entangled and the one that is outside the horizon is emitted as black hole radiation with the other one inside the horizon falling into the black hole. Thus, the entanglement between the emitted radiation and the black hole interior is growing (i.e. in Fig.2 the region of the Cauchy slices behind the black hole horizon is growing exponentially with the time evolution). This monotonic growth superficially corresponds to the size of the black hole interior which is expanding with time like an expanding universe. This suggests that there is no violation of quantum mechanics at each instant. The trouble is caused by the Bekenstein-Hawking entropy formula which states that the system, i.e. the black hole, that is entangled with the emitted radiation $R$ has a finite Hilbert space.
@@ 3 Global Symmetry in the island setup
We will consider a slightly simpler setup than the one in Sec. 2. We consider gravitational empty AdS coupled to a bath. Ref.Almheiri:2019psf ; Penington:2019npb ; Geng:2020qvw showed that entanglement islands exist in such a setup. We consider an empty AdSd geometry, which has a dynamical U(1) gauge field $A_{\mu}(x,z)$ with a charged complex scalar field $\phi(x,z)$ having charge $q$ that leaks to a non-gravitational bath (see Fig.8). We take the non-gravitational bath to be a $d$-dimensional conformal field theory with an a priori independent U(1) global symmetry that lives on half-Minkowski space. We couple the AdSd and the bath such that the U(1) charge in the AdS can leak into the bath. [^3] We emphasize that the AdS is gravitational but since we don’t explicitly study gravitational dynamics in this paper besides taking into account of the existence of the island and using the AdS/CFT dictionary, we focus on the gauge field and charged matter sector. The explicit descriptions of this setup are presented in the following subsections from both the dual field theory perspective and the point of view of the island setup itself.
@@ 3.1 The Boundary Field Theory Analysis
In the dual island setup, the gravitational AdS contains a gauge field $A_{a}(x,z)$ whose holographic dual is $J_{1}^{\mu}(x)$. From Equ. (6) we can see that the bath coupling breaks this U(1) symmetry. Thus the current operator $J_{1}^{\mu}(x)$ is not protected from picking up an anomalous dimension. According to the AdS/CFT correspondence, this anomalous dimension will be dual to the mass of the gauge field $A_{a}(x,z)$. The leading order $g^{2}$ contribution to the anomalous dimension can be computed using the conformal algebra combined with the operator/state correspondence.Aharony:2006hz ; Karch:2023wui . The calculation goes as following.
@@ 3.2.1 The Gauge Symmetry is Spontaneously Broken
The complex scalar field $\phi_{1}(x,z)$ obeys the transparent boundary condition near the asymptotic boundary $z\rightarrow 0$, so the charge it carries leaks to the bath according to Equ. (5). The coupling terms in Equ. (5) can be thought of as double-trace deformations of the standard AdS/CFT setup. According to Witten:2001ua ; Geng:2023ynk , we have the near boundary expansion of the bulk scalar field $\phi_{1}(x,z)$
@@
To show that the U(1) gauge symmetry is spontaneously broken and compute the gauge boson mass, let’s restore the full U(1) gauge invariance by introducing a compensating field $\theta(x,z)$ PRESKILL1991323 which transforms nonlinearly under the gauge transform as
@@
which obey $g_{ab^{\prime}}n^{b^{\prime}}=-n_{a}$ Allen:1985wd . Then, the current-current two-point function Equ. (39) can be decomposed as
@@
We can get the dual boundary $C_{J}$ from the bulk gauge field two-point function following D'Hoker:1999jc as

### v2

@@ Abstract
A central conjecture in quantum gravity is the non-existence of global symmetries. As a fully unitary theory, there is no information loss in a UV complete quantum gravity theory. We see both these concepts reflected in the AdS/CFT correspondence, which tells us that dynamical processes in AdS are fully captured by a manifestly unitary CFT with no information loss. Furthermore, global symmetries of the CFT are dual to gauge symmetries in the AdS, which implies no global symmetry in the AdS. In this work, we provide concrete evidence for the connection between the non-existence of global symmetries and the absence of information loss in quantum gravity. We study the island setups in which a gravitational AdS is coupled with a nongravitational bath on its boundary. In such theories, the information in the AdS can be lost to the bath. We provide concrete examples with global symmetries in the island setup, from both the bottom-up and the top-down perspectives. We argue that these global symmetries are consistent due to entanglement islands, in which holography is realized in a novel fashion. The global symmetries we construct are all mixed with spontaneously broken gauge symmetries. We will show that this fact has two implications: 1) The black hole hair is detectable in the bath (“wet hair”); 2) a resolution of a puzzle proposed by Harlow and Shaghoulian.
@@ 1 Introduction
It has long been conjectured that global symmetries cannot exist in a UV complete quantum gravity theory Banks and Dixon (1988); Giddings and Strominger (1988); Dine (1992); Kallosh et al. (1995); Arkani-Hamed et al. (2007); Banks and Seiberg (2011) and therefore symmetries should be either gauged or broken. This conjecture about the absence of global symmetry applies for both continuous and discrete global symmetries. In this paper, we focus on continuous global symmetries. The basic argument for the absence of this type of global symmetries makes use of the Bekenstein-Hawking entropy formula for a black hole
@@
and the no-hair theorem of black holes. The Bekenstein-Hawking formula suggests that a black hole has a finite number of microstates upper bounded by $e^{S_{BH}}$ and the no-hair theorem states that black hole solutions are independent of the global symmetry charge inside the black hole. The latter suggests that coarse-grained properties of black holes are not affected by adding any amount of global symmetry charges, which however is in tension with the former as it suggests that there would exist an infinite number of black hole microstates distinguished by their global symmetry charges. Hence, global symmetries had better not exist in quantum gravity. We note that within AdS/CFT, a putative proof of the absence of global symmetries has been put forward in Harlow and Ooguri (2019, 2018) (see also Chen and Lin (2021); Hsin et al. (2021); Bah et al. (2023) for some other attempts from a rather different perspective).
@@
The finite Hilbert space the logarithm of whose dimension scales linearly with the black hole horizon area is the early motivation for the holographic principle ’t Hooft (1993); Susskind (1995). This principle is explicitly realized by the AdS/CFT correspondence Maldacena (1999); Gubser et al. (1998); Witten (1998), which states that quantum gravitational dynamics in an asymptotically anti-de Sitter (AdS) spacetime is dual to a conformal field theory (CFT) living on its asymptotic boundary. In the context of a black hole in AdS space, the quantum mechanical system that describes its physics is the dual boundary CFT, which is manifestly unitary. Hence, the Bekenstein-Hawking entropy formula suggests that the black hole dynamics, or the quantum gravitational theory that describes the black hole, is unitary. We also have seen from the previous paragraph that the Bekenstein-Hawking entropy formula suggests the absence of global symmetries in quantum gravity. Therefore, it is natural to conjecture that the absence of global symmetry is tied to the unitarity of quantum gravity.
@@
In this paper, we provide concrete evidence for the above conjecture from a novel perspective. We study a system in which a gravitational asymptotically AdS spacetime is coupled to a nongravitational bath on its asymptotic boundary, so that the gravitational theory is no longer unitary.  [^1] In principle, one can violate unitarity for a finite time, in which case our arguments still apply but the global symmetry can exist only increasingly deeper in the bulk. Such setups, with a coupling to a bath, contain entanglement islands, which have been important to recent developments in quantum gravity Almheiri et al. (2019); Penington (2020); Geng and Karch (2020); Chen et al. (2020a, b), so we call them island setups for convenience. Notice that our definition of island setups does not include all theories with islands, but only those in which unitarity is violated by the coupling of the gravitational system to a bath. The so-called Page curve of the black hole radiation can be computed in this setup, which suggests that the dynamics of the union of the gravitational AdS and the bath is unitary. However, in this system the dynamics in the gravitational AdS alone is not unitary due to the bath coupling. Thus, if global symmetries exist in such systems, it would demonstrate the connection between unitarity and the absence of global symmetries. More precisely, one should expect that particles charged under global symmetries can propagate into the gravitational AdS from the bath. Nevertheless, one might naively think that this expectation violates our former considerations of black holes in AdS. We will show in this paper that this is not the case and the reason is the existence of entanglement islands. We will show how entanglement islands allow for consistency of global symmetries with the above considerations of black holes.
@@
We will furthermore provide explicit examples of global symmetries in the island setup from both bottom-up holographic constructions and top-down string theory constructions. With the general argument that global symmetries can exist in the island setup and various concrete examples, we prove the early conjecture that global symmetries can exist in quantum gravitational systems with disordering Benini et al. (2023). Moreover, our work provides firm evidence for the conjecture that unitarity is essential to the absence of global symmetries in quantum gravity. To be clear we are considering only theories in which the full system preserves unitarity, but the AdS system on its own does not.
@@
We note that Ref. Harlow and Shaghoulian (2021) proposed that unitarity and the absence of global symmetries in gravitational theories can be connected. We attack this question with different considerations and conclusions in specific examples. We should also mention the possibility that this connection might be violated in some low dimensional theories Harlow and Shaghoulian (2021). We do not address this question here but focus on the role of unitarity in quantum gravity theories which have propagating gravitational degrees of freedom.
@@
This paper is organized as follows. In Sec. 2 we review the physics of entanglement islands and provide a general argument of why global symmetry can exist in the island setup. In Sec. 3 we provide an explicit construction of a global symmetry in the island setup. In Sec. 4 we study the holographic dual of the island setup, i.e. the Karch-Randall braneworld Karch and Randall (2001b, c) and we construct global symmetries in the dual island setup using holography. In Sec. 5 we provide constructions of global symmetries for the dual island setup in the string theory realization of the Karch-Randall braneworld. In Sec. 6 we resolve a puzzle proposed by Harlow and Shaghoulian in Harlow and Shaghoulian (2021) where they argued that global symmetries are in tension with entanglement islands. We conclude our paper with discussions in Sec. 7.
@@ 2.1 A Lightening Review of Entanglement Islands
Entanglement islands emerged from attempts to compute the Page curve of the radiation from a black hole. This was a hard question for both technical and conceptual reasons. At the conceptual level, an evaporating black hole’s radiation and the black hole itself are living in the same spacetime, so it is unclear how to isolate the radiation from the black hole. At the technical level, it is unclear how the computation of fine-grained properties like the entanglement entropy can be done in a gravitational spacetime. Thus, a natural setup that can circumvent the above difficulties is one with a black hole spacetime coupled to a non-gravitational bath by a transparent boundary condition for energy on its boundary. Black holes in AdS have a nice asymptotic structure which allows the transparent boundary condition to be clearly modeled Porrati (2002); Porrati and Starinets (2002); Porrati (2003); Duff et al. (2004); Aharony et al. (2006); Geng (2023b) (see Fig.1). In this setup, one can think of the non-gravitational bath as extracting the radiation from the black hole, so one can study the black hole radiation by studying this non-gravitational bath (see Fig.2).
An interesting discovery from studying the above setup is the entanglement island. It comes from the attempt to compute the entanglement entropy of the black hole radiation. The black hole radiation can be thought of as a subregion $R$ in the non-gravitational bath. The entanglement entropy of it can be computed using the quantum extremal surface formula Engelhardt and Wall (2015)
@@
where $\mathcal{I}$ is a subregion in the gravitational AdS black hole spacetime, $S_{\text{QFT}}(R\cup\mathcal{I})$ is the entanglement entropy for the quantum fields in the subregion $R\cup\mathcal{I}$ as if it is in a fixed non-gravitational background, $A(\partial\mathcal{I})$ is the area of the boundary of the subregion $\mathcal{I}$ and $G_{N}$ is the AdS Newton’s constant. In the above formula, the subregion $\mathcal{I}$ is called the entanglement island of $R$. This formula can be derived by using holography to translate the field theory replica path integral for $S(R)$ to a gravitational path integral in the full island setup and the replica wormhole saddle in the resulting gravitational path integral proves Equ. (2) Almheiri et al. (2020); Penington et al. (2019); Geng (2024); Geng et al. (2025). The results of the formula Equ. (2) are the value of the entanglement entropy of the bath subregion $R$ and the location of its entanglement island $\mathcal{I}$ within the gravitational AdS black hole spacetime. The dominance of the entanglement island in the path integral emerges at late time (see Fig.3) and the resulting $S(R)$ obeys a unitary Page curve (see Fig.4).
Physically, one can think that the bath subregion $R$ is capturing more and more radiation from the black hole as time evolves (see Fig.2). Thus, its entanglement entropy is initially growing with time according to Hawking’s calculation Hawking (1975) and this growth is saying that the entanglement between the black hole and the radiation it already emitted is becoming stronger. Nevertheless, a basic consistency from quantum mechanics is that the entanglement entropy between two quantum systems, i.e. the black hole and the radiation $R$, is upper bounded by the smallest Hilbert space dimension among these two quantum systems, i.e.
@@
However, a careful interpretation is required if one compares the above results with Hawking’s calculation Hawking (1975). Hawking’s calculation can be interpreted heuristically as virtual pairs of particles being continuously produced near the black hole horizon, due to the lack of a global time translation invariance extending into that region, and each pair consists of two particles with one outside black hole horizon and one inside (see Fig.5). The two particles in each pair are maximally entangled and the one that is outside the horizon is emitted as black hole radiation with the other one inside the horizon falling into the black hole. Thus, the entanglement between the emitted radiation and the black hole interior is growing (i.e. in Fig.2 the region of the Cauchy slices behind the black hole horizon is growing exponentially with the time evolution). This monotonic growth superficially corresponds to the size of the black hole interior which is expanding with time like an expanding universe. This suggests that there is no violation of quantum mechanics at each instant. The trouble is caused by the Bekenstein-Hawking entropy formula which states that the system, i.e. the black hole, that is entangled with the emitted radiation $R$ has a finite Hilbert space.
@@ 3 Global Symmetry in the island setup
We will consider a slightly simpler setup than the one in Sec. 2. We consider gravitational empty AdS coupled to a bath. Ref.Almheiri et al. (2019); Penington (2020); Geng and Karch (2020) showed that entanglement islands exist in such a setup. We consider an empty AdSd geometry, which has a dynamical U(1) gauge field $A_{\mu}(x,z)$ with a charged complex scalar field $\phi(x,z)$ having charge $q$ that leaks to a non-gravitational bath (see Fig.8). We take the non-gravitational bath to be a $d$-dimensional conformal field theory with an a priori independent U(1) global symmetry that lives on half-Minkowski space. We couple the AdSd and the bath such that the U(1) charge in the AdS can leak into the bath. [^3] We emphasize that the AdS is gravitational but since we don’t explicitly study gravitational dynamics in this paper besides taking into account of the existence of the island and using the AdS/CFT dictionary, we focus on the gauge field and charged matter sector. The explicit descriptions of this setup are presented in the following subsections from both the dual field theory perspective and the point of view of the island setup itself.
@@ 3.1 The Boundary Field Theory Analysis
In the dual island setup, the gravitational AdS contains a gauge field $A_{a}(x,z)$ whose holographic dual is $J_{1}^{\mu}(x)$. From Equ. (6) we can see that the bath coupling breaks this U(1) symmetry. Thus the current operator $J_{1}^{\mu}(x)$ is not protected from picking up an anomalous dimension. According to the AdS/CFT correspondence, this anomalous dimension will be dual to the mass of the gauge field $A_{a}(x,z)$. The leading order $g^{2}$ contribution to the anomalous dimension can be computed using the conformal algebra combined with the operator/state correspondence.Aharony et al. (2006); Karch et al. (2024). The calculation goes as following.
@@ 3.2.1 The Gauge Symmetry is Spontaneously Broken
The complex scalar field $\phi_{1}(x,z)$ obeys the transparent boundary condition near the asymptotic boundary $z\rightarrow 0$, so the charge it carries leaks to the bath according to Equ. (5). The coupling terms in Equ. (5) can be thought of as double-trace deformations of the standard AdS/CFT setup. According to Witten (2001); Geng (2023b), we have the near boundary expansion of the bulk scalar field $\phi_{1}(x,z)$
@@
To show that the U(1) gauge symmetry is spontaneously broken and compute the gauge boson mass, let’s restore the full U(1) gauge invariance by introducing a compensating field $\theta(x,z)$ Preskill (1991) which transforms nonlinearly under the gauge transform as
@@
which obey $g_{ab^{\prime}}n^{b^{\prime}}=-n_{a}$ Allen and Jacobson (1986). Then, the current-current two-point function Equ. (39) can be decomposed as
@@
We can get the dual boundary $C_{J}$ from the bulk gauge field two-point function following D’Hoker et al. (1999) as